Find the percent of the number.
11. 25% of 60
Any help with this equation would be Greatly appreciated.

Answer by jim_thompson5910(28595)

First convert 25% to a decimal number by dividing 25 by 100

25%=25/100=0.25


Now multiply 0.25 by 60 to get

0.25*60=15


So 25% of 60 is 15